Let A = {(x, y, z) E R^3ㅣx - y + z = 0} where E represents "an element of" and R the set of real numbers
(a) Show that A is a vector space
(b) Give the dimensions and basis of A.
1
Expert's answer
2021-02-01T07:31:59-0500

Let "A = \{(x, y, z) \in\mathbb R^3ㅣx - y + z = 0\}"



(a) Let us show that "A" is a vector space. For this it is sufficient to show that "A" is a subspace of the vector space "\mathbb R^3". Let "(x_1,y_1,z_1),(x_2,y_2,z_2)\in A,\ \ \alpha\in \mathbb R". Then "x_1-y_1+z_1=0" and "x_2-y_2+z_2=0". Since "(x_1+x_2)-(y_1+y_2)+(z_1+z_2)=(x_1-y_1+z_1)+(x_2-y_2+z_2)=0+0=0" and "\alpha x_1-\alpha y_1+\alpha z_1=\alpha( x_1-y_1+z_1)=\alpha\cdot 0=0", we conclude that "(x_1,y_1,z_1)+(x_2,y_2,z_2)=(x_1+x_2,y_1+y_2,z_1+z_2)\in A" and "\alpha (x_1,y_1,z_1)=(\alpha x_1,\alpha y_1,\alpha z_1)\in A". Therefore, "A" is a vector space.


(b) Let us give the dimensions and basis of "A". Since "A = \{(x, y, z) \in\mathbb R^3ㅣx - y + z = 0\}= \{(x, x+z, z) \in\mathbb R^3ㅣx, z \in\mathbb R\}=
 \{(x, x, 0)+(0, z, z) \in\mathbb R^3ㅣx, z \in\mathbb R\}=
 \{x(1, 1, 0)+z(0, 1, 1) \in\mathbb R^3ㅣx, z \in\mathbb R\},"

we conclude that the lenearly independent vectors "(1,1,0)" and "(0,1,1)" form a basis of "A". Consequently, the dimension of "A" is equal to 2.